# DUFOMap: Efficient Dynamic Awareness Mapping

### 03 Mar 2024 v1

@@ Abstract
The dynamic nature of the real world is one of the main challenges in robotics. The first step in dealing with it is to detect which parts of the world are dynamic. A typical benchmark task is to create a map that contains only the static part of the world to support, for example, localization and planning. Current solutions are often applied in post-processing, where parameter tuning allows the user to adjust the setting for a specific dataset. In this paper, we propose DUFOMap, a novel dynamic awareness mapping framework designed for efficient online processing. Despite having the same parameter settings for all scenarios, it performs better or is on par with state-of-the-art methods. Ray casting is utilized to identify and classify fully observed empty regions. Since these regions have been observed empty, it follows that anything inside them at another time must be dynamic. Evaluation is carried out in various scenarios, including outdoor environments in KITTI and Argoverse 2, open areas on the KTH campus, and with different sensor types. DUFOMap outperforms the state of the art in terms of accuracy and computational efficiency. The source code, benchmarks, and links to the datasets utilized are provided. See https://kin-zhang.github.io/dufomap for more details.
@@ I Introduction
Several methods are proposed for this. Learning-based methods [3, 4, 5] require training data and often lack explainability. In contrast, methods based on geometric analysis, such as ray casting and visibility [6, 7, 8, 9], often do not support online execution, as they rely on prior maps for difference calculations, and are computationally and memory expensive. Furthermore, they often require parameter tuning for each new setting.
@@ II-B Geometric Analysis Methods
Two of the most popular post-process methods are Removert [7] and ERASOR [8]. They first build a map from all the point clouds and are thus confined to offline operation. Removert [7] projects the map into range images at the location of each query scan. Dynamic points are found based on visibility constraints by comparing query and map range images and using voting. To reduce the number of false positives, range images at decreasing resolution are used to revert dynamic points back to being static.
@@
Occupancy grids, such as OctoMap [14] and UFOMap [10], use ray casting to update the occupancy values of the voxels in a 3D grid. This results in an estimate of the probability of each voxel being occupied over time. Points that fall into cells with an occupancy probability above some threshold are considered static. Both offline and online operation is possible. For online execution, the classification of a point is performed when the point is acquired.
@@ III-B Dealing With the Real World
III-C Classifying Points as Dynamic or Static
IV Experimental Setup
We compare our method with the current state-of-the-art represented by Removert [7], ERASOR [8], OctoMap [14] and Dynablox [15] [^1] All codes of benchmark methods and datasets are available at https://github.com/KTH-RPL/DynamicMap_Benchmark.. The first three are evaluated in post-processing mode against DUFOMap. Dynablox is an online method and is compared to ${}^{\star}$.
IV-A Datasets
To demonstrate that our method can handle a wide range of scenarios and sensor types, we go beyond the use of a single dataset. To achieve this, we follow the benchmark evaluation protocol presented in [9]. The benchmark includes the KITTI dataset [18] with annotation labels and poses from SemanticKITTI [19]. We present results on sequences 00 and 01 in the paper. These comprise a small town and a highway, respectively, and are captured with a HDL-64E LiDAR. For other KITTI sequences, we refer to our project page https://kin-zhang.github.io/dufomap. Another dataset [20] collected by two VLP-32C LiDAR sensors is ‘Argoverse 2 big city’ includes various dynamic objects in an urban environment. The most sparse one is collected by a 16-channel LiDAR in a highly structured (‘Semi-indoor’) environment as illustrated in Fig. 5.
[…]
IV-B Metric
IV-C Parameter Settings
[…]
IV-D Hardware
Experiments were carried out on a desktop equipped with an Intel Core i9-12900KF. To assess real-time robot applicability, we also performed experiments on a robot equipped with an Intel® NUC with an Intel Core i7-8559U.
@@ V-A1 Accuracy
Looking at the online methods, we see that as expected, ${}^{\star}$ performs worse than DUFOMap, which has access to all data before classification is performed. Also, as expected, the performance drop is the smallest in the dataset with the least complex dynamics (KITTI small town). Dynablox does the best on the Argoverse 2 dataset where dynamic objects are constantly moving.
@@ Methods
KITTI highway
Semi-indoor
Removert [7]
0.134 $\pm$ 0.004
0.515 $\pm$ 0.024
ERASOR [8]
0.718 $\pm$ 0.039
0.064 $\pm$ 0.011
@@
Both online methods have low DA for the semi-indoor dataset. One of the two people is standing still at the beginning of the dataset and both are still later (see Fig. 5d). ${}^{\star}$ sees these points as dynamic as soon as the person moves for the first time. Dynablox requires the person to move within a time window to detect it as dynamic.
@@
KITTI [18] SemanticKITTI [19, 23] Kiss-ICP [24]
@@ V-A2 Execution Time
Table II presents information on the run time of the different methods for two of the datasets, one with a 64-channel LiDAR (KITTI highway) and one with a 16-channel LiDAR (semi-indoor). The reported time is the total processing time divided by the number of point clouds. Note that this is mainly important for the three latter methods, for which an online mode is supported. However, we see that OctoMap is prohibitively slow, requiring 3s to integrate a single frame in KITTI.
@@ V-B2 Survey Sensor Dataset
The output of our method is depicted in Fig. 1 (lower right) when applied to the raw data (upper). In the detection results (lower left), we naively clustered the points so that different objects stand out. For a comprehensive view of our results, readers are encouraged to visit our project page [^2] https://kin-zhang.github.io/dufomap where a detailed video of the entire process is available for viewing.
V-C Influence of Pose
In this section, we study how the pose influences the results. We investigate three different sources of pose estimates. The KITTI dataset comes with poses that are used as ground truth in the KITTI odometry benchmark. The poses of the SemanticKITTI dataset are estimated by SuMa [23]. The third method is Kiss-ICP [24]. In Table III we can see that all methods are influenced by the pose. With worse pose estimates, the scenes will appear more dynamic. Therefore, it will be more difficult to classify the static part, and we would expect SA to decrease with increased pose errors. Based on this, Kiss-ICP provides more accurate and consistent pose estimates for the short sequences used in the benchmark, closely followed by SemanticKITTI.
@@ Acknowledgment
Thanks to HKUST Ramlab’s members: Bowen Yang, Lu Gan, Mingkai Tang, and Yingbing Chen, who help collect additional datasets. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.

### 12 Apr 2024 v2

@@ Abstract
The dynamic nature of the real world is one of the main challenges in robotics. The first step in dealing with it is to detect which parts of the world are dynamic. A typical benchmark task is to create a map that contains only the static part of the world to support, for example, localization and planning. Current solutions are often applied in post-processing, where parameter tuning allows the user to adjust the setting for a specific dataset. In this paper, we propose DUFOMap, a novel dynamic awareness mapping framework designed for efficient online processing. Despite having the same parameter settings for all scenarios, it performs better or is on par with state-of-the-art methods. Ray casting is utilized to identify and classify fully observed empty regions. Since these regions have been observed empty, it follows that anything inside them at another time must be dynamic. Evaluation is carried out in various scenarios, including outdoor environments in KITTI and Argoverse 2, open areas on the KTH campus, and with different sensor types. DUFOMap outperforms the state of the art in terms of accuracy and computational efficiency. (See https://kth-rpl.github.io/dufomap for more details.)
Index Terms:
Mapping; Object Detection, Segmentation and Categorization; Robotics and Automation in Construction
@@ I Introduction
Several methods are proposed for this. Learning-based methods [3, 4, 5, 6] require training data and often lack explainability. In contrast, methods based on geometric analysis, such as ray casting and visibility [7, 8, 9, 10], often do not support online execution, as they rely on prior maps for difference calculations, and are computationally and memory expensive. Furthermore, they often require parameter tuning for each new setting.
@@ II-B Geometric Analysis Methods
Two of the most popular post-process methods are Removert [8] and ERASOR [9] with its follow-up ERASOR2 [15]. They first build a map from all the point clouds and are thus confined to offline operation. Removert [8] projects the map into range images at the location of each query scan. Dynamic points are found based on visibility constraints by comparing query and map range images and using voting. To reduce the number of false positives, range images at decreasing resolution are used to revert dynamic points back to being static.
@@
Occupancy grids, such as OctoMap [16] and UFOMap [11], use ray casting to update the occupancy values of the voxels in a 3D grid. This results in an estimate of the probability of each voxel being occupied over time. Points that fall into cells with an occupancy probability above some threshold are considered static. Both offline and online operations are possible. For online execution, the classification of a point is performed when the point is acquired.
@@ III-B Dealing With the Real World
III-C Classifying Points as Dynamic or Static
IV Experimental Setup
We compare our method with the current state-of-the-art represented by Removert [8], ERASOR [9], OctoMap [16] and Dynablox [17]. The first three are evaluated in post-processing mode against DUFOMap. Dynablox is an online method and is compared to ${}^{\star}$.
IV-A Datasets
To demonstrate that our method can handle a wide range of scenarios and sensor types, we go beyond the use of a dataset with a single sensor type. To achieve this, we follow the benchmark evaluation protocol presented in [10]. The benchmark includes the KITTI dataset [20] with annotation labels and poses from SemanticKITTI [21]. We present results on sequences 00 and 01 in the paper. These comprise a small town and a highway, respectively, and are captured with a HDL-64E LiDAR. For other KITTI sequences, we refer to our project page https://kth-rpl.github.io/dufomap. Another dataset [22] collected by two VLP-32C LiDAR sensors is ‘Argoverse 2 big city’ includes various dynamic objects in an urban environment. The most sparse one is collected by a 16-channel LiDAR in a highly structured (‘Semi-indoor’) environment as illustrated in Fig. 5.
[…]
IV-B Metric
IV-C Parameter Settings
[…]
IV-D Hardware
Experiments were carried out on a desktop equipped with an Intel Core i9-12900KF. To assess real-time robot applicability, we also performed experiments on a robot equipped with an Intel NUC with an Intel Core i7-8559U.
@@ V-A1 Accuracy
Looking at the online methods, we see that, as expected, ${}^{\star}$ performs worse than DUFOMap, which has access to all data before classification is performed. Also, as expected, the performance drop is the smallest in the dataset with the least complex dynamics (KITTI small town). Dynablox does the best on the Argoverse 2 dataset where dynamic objects are constantly moving.
@@
KITTI highway Semi-indoor
Removert [8] 0.134 $\pm$ 0.004 0.515 $\pm$ 0.024
ERASOR [9] 0.718 $\pm$ 0.039 0.064 $\pm$ 0.011
@@
Both online methods have low DA for the semi-indoor dataset. One of the two people is standing still at the beginning of the dataset and both are still later (see Fig. 5(d)). ${}^{\star}$ sees these points as dynamic as soon as the person moves for the first time. Dynablox requires the person to move within a time window to detect it as dynamic.
@@
GT poses of KITTI [20] SuMa [21, 25] KISS-ICP [26]
@@ V-A2 Execution Time
[…]
@@ V-B2 Survey Sensor Dataset
The output of our method is depicted in Fig. 1 (lower right) when applied to the raw data (upper). In the detection results (lower left), we naively clustered the points so that different objects stand out. For a comprehensive view of our results, readers are encouraged to visit our project page [^1] https://kth-rpl.github.io/dufomap where a detailed video of the entire process is available for viewing.
V-C Influence of Pose Estimation
In this section, we study how the pose influences the results. We investigate three different sources of pose estimates. The KITTI dataset comes with poses that are used as ground truth in the KITTI odometry benchmark. The ground truth poses of the SemanticKITTI dataset are estimated by SuMa [25]. The third method is KISS-ICP [26]. In Table III we can see that all methods are influenced by the pose. With worse pose estimates, the scenes will appear more dynamic. Therefore, it will be more difficult to classify the static part, and we would expect SA to decrease with increased pose errors. Based on this, KISS-ICP provides more accurate and consistent pose estimates for the short sequences used in the benchmark, closely followed by the ground truth pose of SemanticKITTI (SuMa).
@@ Acknowledgment
Thanks to HKUST Ramlab’s members: Bowen Yang, Lu Gan, Mingkai Tang, and Yingbing Chen, who help collect additional datasets. We also thank the anonymous reviewers for their constructive comments.